# E-Trojans: Ransomware, Tracking, DoS, and Data Leaks on Battery-powered Embedded Systems

### v1

@@ 4 E-Trojans Attacks
We present E-Trojans, four novel attacks on the internals of Xiaomi e-scooters. For example, they can undervolt and destroy the battery, track users by advertising their fingerprint over BLE, and exfiltrate confidential data. Each attack starts by flashing a malicious BCTRL firmware on the victim’s e-scooter and turning OFF legitimate firmware updates (only on the BCTRL). This firmware contains custom code enabling the malicious capabilities required to conduct our attacks.
@@ 4.1 Undervoltage Battery Ransomware (UBR)
UBR is a novel undervoltage ransomware that progressively and irreversibly damages the e-scooter’s battery until the victim pays a ransom. The damage is caused by flashing a malicious firmware that keeps the battery cell voltages below 2750mV or 1580mV (i.e., battery level below 0%). As a result, the battery longevity is reduced [61, 27], and, by keeping the battery undervolted and imbalanced, the attacker can cause short circuits [23].
In the optimal scenario, the attacker undervolts a battery cell as low as 0V and quickly destroys it. We highlight that our attack can damage the battery without asking for a ransom. Replacing a damaged battery can cost up to USD 139 [74]. UBR has five steps, also shown in Algorithm 2:
@@
    The attacker flashes a malicious BCTRL firmware containing the ransomware that exploits remote and proximity-based techniques presented in [12]. The ransomware remains stealthy, letting the user drive as usual. The user rides the e-scooter, discharging it until the battery is low. Although UBR does not require a specific battery level, batteries closer to 0% are faster to undervolt. Then, the user charges the battery using the stock charger.
@@
    UBR enters its initialization phase (first five lines of Algorithm 2), where it: (1) disables undervoltage protections by increasing dUVT to 1580mV and bypassing cUVT safety checks; (2) disables load balancing by bypassing the dLBD and cLBD safety checks; (3) disables battery charging by setting the canCharge flag to False; (4) activates the anti-theft software lock to prevent the e-scooter from becoming inactive and automatically turning OFF; (5) enforces faster discharging by disabling sleep mode and, optionally, spoofing a UART command that turns ON the headlight and taillight.
@@
    The victim installs the ransom app and can monitor the degradation of the battery. They stop the attack by transferring money to the attacker’s crypto wallet. Once the app’s backend confirms the payment (made in cryptocurrency, which complicates tracing), it reveals the secret unlock code, unique to that infected device, to the app’s frontend. The app then sends a special BLE packet that contains this unlock code, activating firmware updates on the BCTRL.
@@
    The ransom app guides the victim through the recovery procedure. The app flashes on the BCTRL, a recovery firmware that allows charging even when the battery is critically undervolted. The stock BCTRL firmware does not allow this for safety reasons (e.g., risk of short circuit). As soon as the battery is not critically undervolted anymore (i.e., voltage is higher than cUVT), the app allows the user to flash the stock BCTRL firmware.
@@ 4.3 Denial of E-Scooter Services (DES)
DES5 disables battery charging by setting the canCharge flag in the BCTRL EEPROM to False. DES6 exploits the error handling routines by impersonating BTS, DRV, or BMON and raising fake errors that put the e-scooter in an irregular state (e.g., alarm noises, locked motor, unresponsiveness, and forced power OFF). DES7 is a denial-of-sleep attack that keeps the BCTRL in a resource-intensive mode instead of letting it switch to sleep mode by altering the BCTRL main loop (described in Algorithm 1).
@@ 4.5 Mapping Attacks and Vulnerabilities
Table 2: Mapping between the four attacks and the four vulnerabilities. All attacks exploit V1, V2, and V3.
Only DES exploits V4.
@@
5 Implementation
5.1 BCTRL Fw Patching and Capabilities
Our Python3 script (payload-patcher.py) binary patches a BCTRL using STM8 assembly code. During this process, the user can include up to ten new capabilities unavailable in the stock BCTRL firmware in the patch. We devised these capabilities thanks to our extensive RE of the BCTRL firmware. The script writes the STM8 assembly code for unused parts of the firmware and then updates the firmware CRC. The resulting binary can be flashed over BLE on the BCTRL of a M365 or ES3 e-scooter. We developed ten capabilities that we describe next.
DFU: Disable Firmware Updates. DFU disables BCTRL firmware updates from Mi Home or third-party apps. We introduce a new canFwUpdate flag (address 0x04CB), set to False by default. We rewrite the HandleUART() function (address 0x94C5) to accept or reject firmware update packets (starting with 0x07, 0x08, 0x09, or 0x0A), depending on canFwUpdate. This flag is set to True only when the BCTRL receives our new unlock packet (starting with 0xEE and followed by the correct unlock code). For example, in UBR, the user retrieves the unique unlock code by paying a ransom and utilizes it to enable firmware updates on the BCTRL.
DDT: Disable Dangerous Thresholds. DDT alters the default BMON undervoltage and overvoltage thresholds. At BCTRL boot, DDT initializes the BMON UV_TRIP (address 0x923D) and OV_TRIP (address 0x923D) registers to 1580mV and 4700mV, their lowest and highest values, instead of the default ones. DDT enables UBR to reach voltages below the undervoltage threshold (i.e., 2750mV) recommended by Xiaomi, resulting in permanent battery damage and increased safety risks.
[…]
SCV: Spoof Cell Voltage. SCV spoofs voltage measurements of battery cells. We rewrite the BCTRL main loop to ignore new BMON measurements and only read the spoofed values we inject in register X. SCV sends the spoofed voltages to the BTS, which are displayed to the user via the Mi Home UI. By displaying fake cell voltage measurements instead of real ones, SCV enables UBR to conceal an undervolted battery from users who check the battery status via Mi Home.
Table 3: Mapping between the ten capabilities and the four attacks. All capabilities enable one or more attacks.
All ten capabilities are required to deploy UBR.
@@ 5.1 BCTRL Fw Patching and Capabilities
5.2 UBR App and Backend
[…]
6 Evaluation
6.1 Setup
We ran our payload-patcher.py script to generate the malicious BMS firmware. We select one of our four attacks in the interactive shell of the script and the target e-scooter model. Then, we set up each attack by connecting to the e-scooter, gaining authorized access (i.e., through malicious pairing or session downgrade) and performing a rogue BCTRL firmware update.
6.2 Results
Attack M365 ES3 Details
UBR  ✓ Battery -50% in 3.5h (M365)
UTI Persists factory resets
DES Unresponsive due to internal failure
PLR Leaks and cracks common passwords
Similarly, we reduced the battery autonomy of the ES3 by around 10% in ten hours, spending six of them depleting the battery to 0% (three hours for the M365). We could not reach a critical undervoltage due to the DRV independently measuring an abnormal voltage and raising Error 24 - Supply Voltage out of range. As such, the ES3 is better protected than the M365 against UBR, even though it still sustained significant damage. We also confirmed that our recovery BCTRL firmware rebalances undervolted batteries in five minutes.
DES. DES made the e-scooter unresponsive and not drivable. DES1 and DES3 cause the DRV to raise Error 21 - No Communication with BMS, halting the motor and powering OFF the e-scooter after ten seconds. DES2 cuts power from all internal systems, powering OFF the e-scooter and barring it from powering ON again. DES4 prevents battery charging regardless of the battery charger hardware or the e-scooter status (e.g., powered ON or OFF). DES5 induces an endless lock state or reboot cycle, rendering the e-scooter unable to move and forcing it continuously to stay powered ON.
PLR. PLR leaked the e-scooter password hash and enabled offline brute forcing. We sniffed three rotating BLE advertisements broadcasted by an infected e-scooter, each containing a fragment of the hash. Since the e-scooter resets after changing its advertisement, we programmed the malicious firmware to rotate it in rapid succession, tricking the user into thinking a single reboot happened instead of three. We also confirmed that retrieving the e-scooter password from the hash is feasible. Our non-optimized rainbow table cracked the seventy passwords, following the most common six-digit patterns, in less than one second on average.
7 Countermeasures
We propose four practical and backward-compliant countermeasures, each addressing one of the four vulnerabilities described in Section 3.3 (i.e., C1 addresses V1, and so on). To fix UBR, UTI, and PLR, Xiaomi must implement C1, C2, and C3. To fix all DES variants, Xiaomi should also add C4.
C1: Encrypt the BCTRL firmware with TEA. Xiaomi should encrypt the BCTRL firmware to protect its confidentiality at rest (e.g., in Mi Home) and in transit (e.g., during a firmware update over BLE). We recommend reusing the TEA block cipher already used to encrypt the DRV firmware. Its overhead is negligible, as TEA is a lightweight cipher [71].
C2: Sign and verify the BCTRL firmware with ECDSA. Xiaomi should digitally sign and verify the BCTRL firmware to prevent rogue BCTRL firmware updates. We recommend reusing ECDSA, which is FIPS-compliant [42] and already employed to sign/verify the DRV and BTS firmware. Its overhead is minimal, as the BTS would have to verify the BCTRL firmware during an update using crypto primitives it already supports.
C3: Protect the UART bus with SCP03. Xiaomi should protect the confidentiality, integrity, and authenticity of the UART bus. We suggest using standard secure embedded bus protocols such as Secure Channel Protocol 03 (SCP03) [45]. SCP03 provides confidentiality, integrity, authenticity, and replay protection using a lightweight scheme based on pre-shared symmetric keys, authenticated encryption, and AES. Since SCP03 was designed for smart cards [43], way more constrained than e-scooters, implementing it adds minimal overhead.
C4: Protect the UART bus with rate limiting. Xiaomi should protect the UART bus against DoS attacks or at least mitigate them. We recommend implementing a lightweight rate-limiting mechanism for UART, such as the leaky bucket algorithm, where excess packets are discarded if the incoming packet rate exceeds the outgoing rate. C4 can be combined with C2 to achieve better DoS protection as an authenticated UART bus discards messages that do not pass the integrity check.
8 Related Work
E-Scooters Security and Privacy. Research on e-scooters has studied their proprietary protocols and rental ecosystems, but no work has focused on their internals. A recent contribution [12] analyzed the Xiaomi e-scooter ecosystem, identifying issues in the proprietary BLE protocols used by the e-scooter and Mi Home. The researchers identified vulnerabilities [1] in the Bird e-scooter sharing platform and discovered MitM attacks [40] on the rental e-scooters of the Lime company. Other studies focused on the privacy of user-related data in e-scooter rental Android apps [68] and the safety risks of riding e-scooters [35].
E-Scooters Hacking and Modding. In 2019, Zimperium revealed flaws in the M365 locking system that could halt a running e-scooter [77]. The same year, the hacker Lanrat found that M365 does not enforce authentication over BLE [21]. ScooterHacking [57], the largest e-scooter modding community, released tools [55, 56] to interact with Xiaomi e-scooters, mod DRV firmware [54] and flash it [7]. In our work, we develop custom BCTRL firmware with capabilities not offered by previous tools and research.
Attacks on Battery. The threats concerning battery overheating, overcharging, and mechanical damage in electric vehicles have been investigated in [18, 31], where the authors discuss thermal runaway. Researchers have proposed a battery drain attack while the ignition is turned OFF [14], an electromagnetic interference attack [66], and a physical access rogue firmware update to a Tesla BMS [29]. Theoretical vulnerabilities in MacBook laptop batteries, capable of overheating or turning it OFF, have been discussed in [38]. The E-Trojans attacks are orthogonal as they target an e-scooter’s battery and BMS.
Attacks on Embedded Firmware. Several studies focused on malicious embedded firmware, but none exploited e-scooter internals. Researchers attacked embedded firmware on programmable control logic (PCL) [22], object trackers [49], printers [17], mice [37], OS management systems [24], mobile bootloaders [46], network equipment [16], botnets [2] and recently industrial wrenches [44].
Attacks on BES. Researchers have explored the security and privacy of various BESs. For example, in [30], they compromised the insecure CAN bus of an automobile by assuming a compromised electric control unit (ECU). Several attacks, such as MitM, DoS, side-channel, privacy leakage, and side-channel, were also found in Wireless charging stations for smartphones [73, 32, 41] and electric vehicle charging infrastructures [50, 3]. Research on drones explored their external attack surfaces, finding vulnerabilities in their firmware update protocol [51], mobile app [34], and the wireless communication channel [9, 20].
9 Conclusion
We develop E-Trojans, a toolkit that implements our attacks by leveraging our RE findings (e.g., disable safety thresholds). We successfully carried out the attacks on actual devices, proving their practicality. For instance, UBR irreversibly reduced the M365 battery autonomy by 50% in three hours and the ES3 battery autonomy by 10% in ten hours. We fix attacks and their root causes with four backward-compliant countermeasures.
Acknowledgments

### v2

@@ 4 E-Trojans Attacks
We present E-Trojans, four novel attacks on the internals of Xiaomi e-scooters. The attacks, among others, can undervolt the e-scooter battery while asking for a ransom, track the e-scooter using internal fingerprints over BLE, and exfiltrate confidential data. Each attack begins by flashing a malicious BCTRL firmware, which includes custom code that enables the necessary malicious capabilities, such as preventing legitimate BCTRL firmware updates. Since firmware updates require authentication, the attacker can bypass this restriction using, for example, the E-Spoofer toolkit [11].
@@ 4.1 Undervoltage Battery Ransomware (UBR)
UBR is a novel undervoltage ransomware that progressively and irreversibly damages the e-scooter’s battery until the victim pays a ransom. The damage is caused by flashing a malicious firmware that keeps the battery cell voltages below 2750mV (i.e., battery level lower than 0%). UBR has five steps, also shown in Algorithm 2:
@@
    The attacker flashes a malicious BCTRL firmware containing the ransomware that exploits remote and proximity-based techniques presented in [12]. The ransomware remains stealthy, letting the user ride the e-scooter as usual, until the battery is low. Although UBR does not require a specific battery level, batteries closer to 0% are faster to undervolt. Then, the user charges the battery using the stock charger.
@@
    UBR enters its initialization phase (first five lines of Algorithm 2), where it: (1) disables undervoltage protections by increasing dUVT to 1580mV and bypassing cUVT safety checks; (2) disables load balancing by bypassing the cLBD safety checks; (3) disables battery charging by setting the canCharge flag to False; (4) activates the anti-theft software lock to prevent the e-scooter from becoming inactive and automatically turning off; (5) enforces faster discharging by disabling sleep mode and, optionally, spoofing a UART command that turns on the headlight and taillight.
@@
    The victim installs the ransom app and can monitor the degradation of the battery. They stop the attack by transferring money to the attacker’s crypto wallet. Once the app’s backend confirms the payment (made in cryptocurrency, which complicates tracing), it reveals the secret unlock code, unique to that compromised device, to the app’s front end. The app then sends a special BLE packet that contains this unlock code, activating firmware updates on the BCTRL.
@@
    The ransom app guides the victim through a recovery procedure to restore the stock BCTRL firmware. The app flashes a BCTRL recovery firmware that allows charging even when the battery is critically undervolted. After the e-scooter is sufficiently charged (i.e., voltage is higher than cUVT), the app allows the user to flash the stock BCTRL.
UBR primarily impacts the victim’s safety by causing battery undervoltage and imbalance, which increases the likelihood of malfunctions and short circuits [28]. In the optimal scenario, the attacker undervolts a battery cell as low as 0mV, reducing its longevity [67, 32]. The impact of UBR, and battery ransomware attacks in general, is high: it causes irreversible physical damage to the battery, which is the most expensive component of an e-scooter. The same reasoning holds for other battery-powered systems, including electric vehicles. Replacing the battery can cost up to USD 139 [81], while most Xiaomi e-scooters are priced between USD 200 and 450. We highlight that UBR can be deployed without asking for a ransom. This variant is a proximity or remote DoS attack that undervolts and destroys the battery without alerting the user.
@@ 4.3 Denial of E-Scooter Services (DES)
DES5 disables battery charging by setting the canCharge flag in the BCTRL’s Flash memory to False. DES6 exploits the error handling routines by impersonating BTS, DRV, or BMON and raising fake errors that put the e-scooter in an irregular state (e.g., alarm noises, locked motor, unresponsiveness, and forced power off). DES7 is a denial-of-sleep attack that keeps the BCTRL in a resource-intensive mode instead of letting it switch to sleep mode by altering the BCTRL main loop (described in Algorithm 1).
@@ 4.5 Mapping Attacks and Vulnerabilities
5 Implementation
5.1 BCTRL Firmware Patch and Capabilities
Our Python3 script (payload-patcher.py) binary patches a BCTRL using STM8 assembly code. During this process, the user can include up to ten new capabilities unavailable in the stock BCTRL firmware in the patch. We devised these capabilities thanks to our extensive RE of the BCTRL firmware. The script writes the STM8 assembly code for unused parts of the firmware and then updates the firmware CRC. The resulting binary can be flashed over BLE on the BCTRL of a M365 or ES3 e-scooter. We developed ten capabilities that now we describe.
DFU: Disable Firmware Updates. DFU disables firmware updates on the BCTRL. We introduce a new canFwUpdate flag (address 0x04CB), set to False by default. We rewrite the HandleUART() function (address 0x94C5) to accept or reject firmware update packets (starting with 0x07, 0x08, 0x09, or 0x0A), depending on canFwUpdate. This flag is set to True only when the BCTRL receives our new unlock packet (starting with 0xEE and followed by the correct unlock code). For example, in UBR, the user retrieves the unique unlock code by paying a ransom and utilizes it to enable firmware updates on the BCTRL.
DDT: Disable Dangerous Thresholds. DDT alters the default BMON undervoltage and overvoltage thresholds. At BCTRL boot, DDT initializes the BMON’s UV_TRIP (address 0x923D) and OV_TRIP (address 0x923D) registers to 1580mV and 4700mV, their lowest and highest values, instead of the default ones. DDT enables UBR to reach voltages below the undervoltage threshold (i.e., 2750mV) recommended by Xiaomi, resulting in permanent battery damage and increased safety risks.
[…]
SCV: Spoof Cell Voltage. SCV spoofs voltage measurements of battery cells. We rewrite the BCTRL main loop to ignore BMON’s voltage measurements and only read the spoofed values we inject in register X. SCV delivers these values to the BTS, displaying them to the user via the Mi Home UI. By faking voltage measurements, SCV allows UBR to hide the real battery voltage of a spoofed cell from users who want to check the battery status via Mi Home.
5.2 UBR App and Backend
[…]
6 Evaluation
6.1 Setup
Our payload-patcher.py script offers an interactive shell to craft malicious BCTRL firmware. The user must select one of our four attacks and a supported e-scooter model (i.e., M365, Pro1, Essential, Pro2, and ES3). Then, we connect to the victim’s e-scooter, gain authorized access (i.e., through malicious pairing or session downgrade) and perform a rogue BCTRL firmware update.
6.2 Results
Attack M365 ES3 Details
UBR  ✓ Battery autonomy -50% in 3.5h (M365)
UTI Persists factory resets
DES Internal component failure
PLR Leaks common passwords
Similarly, we reduced the battery autonomy of the ES3 by around 10% in ten hours, spending six of them depleting the battery to 0% (three hours for the M365). We could not reach a critical undervoltage due to the DRV independently measuring an abnormal voltage and raising Error 24 - Supply Voltage out of range. As such, the ES3 is better protected than the M365 against UBR, even though it still sustained significant damage.
DES. DES made the e-scooter unresponsive and not rideable. DES1 and DES3 cause the DRV to raise Error 21 - No Communication with BMS, halting the motor and powering off the e-scooter after ten seconds. DES2 cuts power from all internal systems, powering off the e-scooter and barring it from powering on again. DES4 prevents battery charging regardless of the battery charger hardware or the e-scooter status (e.g., powered on or off). DES5 induces an endless lock state or reboot cycle, rendering the e-scooter unable to move and forcing it continuously to stay powered on.
PLR. PLR leaked the e-scooter password hash and enabled offline brute forcing. We sniffed three rotating BLE advertisements broadcasted by a compromised e-scooter, each containing a fragment of the hash. Since the e-scooter resets after changing its advertisement, we programmed the malicious firmware to rotate it in rapid succession, tricking the user into thinking a single reboot happened instead of three. We also confirmed that retrieving the e-scooter password from the hash is feasible. Our non-optimized rainbow table cracked the seventy passwords, following the most common six-digit patterns, in less than one second on average.
Safety Concerns. Our evaluation confirmed that a malicious BCTRL can turn off hardware protections in the BCTRL, e.g., load balancing, and in the BMON, e.g., voltage thresholds.
The risk of fires and thermal runaway due to undervoltage is realistic. For example, UBR undervolted the M365’s battery below 100mV (below cUVT = 1580mV), placing it in a harmful deep discharge state that increases the likelihood of short circuits and fires. This dangerous condition is also difficult for users to detect, as the BCTRL can spoof battery parameters to conceal the damage.
Battery overheating via overvoltage requires an additional assumption, i.e., a compromised or malfunctioning charger. We confirmed the feasibility of overvoltage using a programmable power supply in a safe environment. We raised the battery voltage to 4900mV (above cOVT = 4700mV), and terminated the experiment to avoid safety risks.
7 Countermeasures
We propose four practical and backward-compatible countermeasures, each addressing one of the four vulnerabilities described in Section 3.3 (i.e., C1 addresses V1, and so on). To fix UBR, UTI, and PLR, Xiaomi must implement C1, C2, and C3. To fix all DES variants, Xiaomi should also add C4.
Our countermeasures introduce negligible overhead, as they leverage lightweight cryptography (TEA [78], SCP03 [50]) designed for embedded systems, while providing valuable security guarantees, including encryption and authentication, and DoS protection. They do not contribute to vendor lock-in for batteries, as third-party batteries remain compatible with the stock BMS.
C1: Encrypt the BCTRL firmware with TEA. Xiaomi should encrypt the BCTRL firmware to protect its confidentiality at rest (e.g., in Mi Home) and in transit (e.g., during a firmware update over BLE). We recommend reusing the TEA block cipher already used to encrypt the DRV firmware.
C2: Sign and verify the BCTRL firmware with ECDSA. Xiaomi should digitally sign and verify the BCTRL firmware to prevent rogue BCTRL firmware updates. We recommend reusing ECDSA, which is FIPS-compliant [48] and already employed to sign/verify the DRV and BTS firmware.
C3: Protect the UART bus with SCP03. Xiaomi should protect the confidentiality, integrity, and authenticity of the UART bus. We suggest using standard secure embedded bus protocols such as Secure Channel Protocol 03 (SCP03) [50]. SCP03 provides confidentiality, integrity, authenticity, and replay protection using a lightweight scheme based on pre-shared symmetric keys, authenticated encryption, and AES.
C4: Protect the UART bus with rate limiting. Xiaomi should protect the UART bus against DoS attacks or at least mitigate them. We recommend implementing an efficient rate-limiting mechanism for UART, such as the leaky bucket algorithm, where excess packets are discarded if the incoming packet rate exceeds the outgoing rate. Combining C4 with C2 enhances DoS protection, as C4 discards messages failing the integrity checks.
8 Related Work
E-Scooters’ Security and Privacy. Research on e-scooters has studied their proprietary protocols and rental ecosystems, but no work has focused on their internals. A recent contribution [12] analyzed the Xiaomi e-scooter ecosystem, identifying issues in the proprietary BLE protocols used by the e-scooter and Mi Home. The researchers identified vulnerabilities [1] in the Bird e-scooter sharing platform and discovered MitM attacks [46] on the rental e-scooters of the Lime company. Other studies focused on the privacy of user-related data in e-scooter rental Android apps [75] and the safety risks of riding e-scooters [40].
E-Scooters Hacking and Modding. In 2019, Zimperium revealed flaws in the M365 locking system that could halt a running e-scooter [83]. The same year, the hacker Lanrat found that M365 does not enforce authentication over BLE [23]. ScooterHacking [63], the largest e-scooter modding community, released tools [61, 62] to interact with Xiaomi e-scooters and perform firmware updates [7]. Modding tools, such as [60], focus on increasing the e-scooter’s performance by, for example, increasing the motor’s speed value stored in the DRV’s Flash memory. In contrast, we focus on the BCTRL firmware and develop custom patches for it. Our patches incorporate a substantial amount of binary instructions, including conditional jumps and new UART packets, that implement new capabilities previously unavailable on a Xiaomi e-scooter.
Attacks on Battery. The likelihood of battery overheating, thermal runaway, overcharging, and mechanical damage in electric vehicles have been investigated in [20, 36]. Researchers have proposed a battery drain attack while the ignition is turned off [15], an electromagnetic interference attack [72], and a physical access rogue firmware update to a Tesla BMS [34]. Theoretical vulnerabilities in MacBook laptop batteries, capable of overheating or turning off the machine, have been discussed in [43]. The E-Trojans attacks are orthogonal as they target the battery and BMS of an e-scooter.
Attacks on Embedded Firmware. Several studies have focused on malicious embedded firmware, but we are the first to introduce a battery ransomware leveraging a malicious BMS firmware. Researchers attacked embedded firmware on programmable control logic (PCL) [24], object trackers [54], printers [19], mice [42], OS management systems [29], mobile bootloaders [51], network equipment [18], botnets [2] and recently industrial wrenches [49].
Attacks on BES. Security research on BESs has identified vulnerabilities across various domains. Automotive researchers exploited the insecure CAN bus by compromising an Electronic Control Unit (ECU) [35]. Similar works have exposed wireless charging stations and EV infrastructure to attacks such as MitM, DoS, and privacy leakage [80, 37, 47, 55, 3]. Studies on drones have revealed issues in firmware updates, mobile apps, and wireless traffic [56, 39, 9, 22]. Several IoT devices, such as fitness trackers [13, 16, 27], vacuum robots [26, 74, 25], crypto wallets [44], and smart doorbells [57], have been found vulnerable to a broad range of attacks, including spoofing, eavesdropping, MitM, privacy, and physical attacks. Compared to prior attacks on BESs, we introduce several novel aspects, including the RE of e-scooters’ internal attack surface (an area previously unexplored in the literature), the development of new proximity and remote attacks (e.g., undervoltage battery ransomware and user tracking), and the evaluation of voltage protections in the Xiaomi BMS, along with their safety implications.
9 Conclusion
We develop E-Trojans, a toolkit that implements our attacks by leveraging our RE findings (e.g., disable safety thresholds). We successfully carried out the attacks on actual devices, proving their practicality. For instance, UBR irreversibly reduced the M365 battery autonomy by 50% in three hours and the ES3 battery autonomy by 10% in ten hours. We fix attacks and their root causes with four backward-compatible countermeasures.
Acknowledgments
Appendix - Additional Tables
Table 2: Battery and BMS details. OVT: overvoltage threshold, UVT: undervoltage threshold, LBD: load balancing delta. Prefix "d" means dangerous, and prefix "c" means critical.
M365 ES3
Battery
Manuf. Ninebot Ninebot
Type Li-Ion 18650 Li-Ion 18650
Chem. 10INR19/66-3 10INR19/66-3
Cells 30 (10x3) 30 (10x3)
Capacity 7.80Ah/300Wh 7.65Ah/275Wh
Voltage 42V to 36V 42V to 36V
BMS
BCTRL STM8L151K6 STM8L151K6
BMON BQ76930 BQ76930
dOVT 4200mV 4200mV
cOVT 4700mV 4700mV
dUVT 2750mV 2750mV
cUVT 1580mV 1580mV
cLBD 800mV 800mV
@@ Acknowledgments
Table 3: Mapping between the four attacks and the four vulnerabilities. All attacks exploit V1, V2, and V3. Only the DES attacks exploit V4.
@@
Table 4: Mapping between the ten capabilities and the four attacks. All capabilities enable one or more attacks. All ten capabilities are required to deploy UBR.